OpenStudy (anonymous):

Solve For X (x-6)(x+10)=0

OpenStudy (anonymous):

to solve for x, we know that at least one of the factors must equal 0 for at least one value of x because anything times 0 is 0 for (x-6) to equal 0, x=6 for (x+10) to equal 0, x=-10 so x= 6,-10